A random variable X has Poisson distribution with mean equal to 0.4. What is the probability that the random variable is greater than zero?
1
Expert's answer
2020-06-01T19:27:06-0400
"P(X=x)={e^{-\mu}\cdot \mu^x \over x!}"

Given "\mu=0.4"


"P(X=0)={e^{-0.4}\cdot 0.4^0 \over 0!}=e^{-0.4}"

"P(X>0)=1-P(X=0)=1-e^{-0.4}\approx0.330"